OpenStudy (anonymous):

Convert 2cis5(pi)/3 3 to rectangular form.

OpenStudy (anonymous):

\[2cis \frac{ 5\pi }{ 3 }\]

OpenStudy (anonymous):

\[-2-2\sqrt{3i}\] Is A

OpenStudy (anonymous):

@dtan5457

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

@skullpatrol

OpenStudy (anonymous):

@amistre64